# Orbits of particles with magnetic dipole moment around magnetized Schwarzschild black holes: Applications to S2 star orbit

### v1

@@ Abstract
This study provides a comprehensive analytical investigation of the bound and unbound motion of magnetized particles orbiting a Schwarzschild black hole immersed in an external asymptotically uniform magnetic field, which includes all conceivable types of bounded and unbounded orbits. In particular, for planetary orbits, we perform a comparative analysis of our findings with the observed position of the S2 star carrying magnetic dipole moment around Sagittarius A* (Sgr A*). We found maximum and minimum values for the parameter of magnetic interaction between the magnetic dipole of the star and the external magnetic field, as well as the energy and angular momentum of the S2 star. As a result, we obtain estimations of the magnetic dipole of the star in order of $10^{6}\rm\ G\cdot cm^{3}$. Additionally, we explore deflecting trajectories akin to gravitational Rutherford scattering. In obtaining the solutions for the orbital equations, we articulate the elliptic integrals and Jacobi elliptic functions, and our study is augmented by illustrative figures and simulations.
@@ I Introduction
The investigation of free-falling objects in gravitational fields links geometry and gravity through the general theory of relativity, where the curvature of spacetime plays the principal role. This fundamental understanding gained popularity early once it was recognized that planets and light follow geodesics. Outstanding discoveries, right after the proposition of the Schwarzschild solution Schwarzschild (1916), was the verification of light deflection during the 1919 solar eclipse expedition Dyson et al. (1920). The precise evaluation of the anomalous precession in the perihelion of Mercury Clemence (1947), which constitute the primary standard tests of general relativity. However, the complex nature of the partial differential equations governing particle dynamics in curved spacetimes necessitates simplified or numerical solutions to compare theoretical results with observational tests.
@@
Analytical expressions have several advantages since they serve as benchmarks for numerical methods, and hence, they facilitate comprehensive parameter space exploration for predicting astrophysical observables. Efforts to find exact analytical solutions for geodesic equations of particles have intensified since Hagihara’s studies in 1931 Hagihara (1930), followed by works by Darwin, Mielnik, and Plebański Darwin (1959, 1961); Mielnik and Plebański (1962). These methods, which are based on the theories of elliptic functions and modular forms, were studied by nominated nineteenth-century mathematicians such as Jacobi Jacobi (2013), Abel Abel , Riemann Riemann (1857, 1866), and Weierstrass Weierstrass (1854) (see also Ref. Baker (1995) for a complete textbook review on these discoveries). Particularly noteworthy is the application of modular forms in solving (hyper-)elliptic integrals involved in particle trajectory studies, garnering significant attention in recent decades. These methods, rooted in the theories of elliptic functions and modular forms, have been the focus of numerous investigations analyzing trajectory curves in black hole spacetimes inferred from general relativity and its extensions (see, for example, Refs. Kraniotis and Whitehouse (2002, 2003); Kraniotis (2004, 2005); Cruz et al. (2005); Kraniotis (2007); Hackmann and Lämmerzahl (2008a, b); Hackmann et al. (2009, 2010); Olivares et al. (2011); Kraniotis (2011); Cruz et al. (2013); Villanueva et al. (2013); Kraniotis (2014); Soroushfar et al. (2015); Villanueva and Olivares (2015); Soroushfar et al. (2016); Hoseini et al. (2016, 2017); Fathi et al. (2020a, b, 2021a); González et al. (2021); Fathi et al. (2021b); Kraniotis (2021); Fathi et al. (2022); Soroushfar and Afrooz (2022); Battista and Esposito (2022); Fathi et al. (2023); Fathi (2023)).
@@
Building upon this scientific foundation, contemporary research continues to explore the properties and dynamics of particles falling in gravitational fields in diverse astrophysical contexts. bound and unbound particle orbits, in particular, represent fundamental aspects of motion that offer valuable insights into the structure and evolution of the universe. Bound orbits, characterized by closed trajectories around central masses, are found extensively in celestial systems, ranging from planetary orbits in the solar system to the motion of stars in galaxies Perek (1962); Barbieri and Gratton (2002). These orbits contribute to astrophysical objects’ stability and long-term dynamics and provide essential constraints on theoretical models of gravitational dynamics and cosmological evolution.
@@
Given the importance of this subject, our study aims to investigate the dynamics of magnetized particles affected by the magnetic field of an astrophysical object, in this case, a magnetized Schwarzschild black hole. In this study, we analyze the dynamics and motion of particles possessing magnetic dipole moments analytically as they move in the exterior geometry of a Schwarzschild black hole with an associated magnetic field. We utilize elliptic integrals to express exact analytical solutions to the equations of motion governing the behavior of these particles. Particularly, the deflecting trajectories are considered akin to a type of gravitational Rutherford scattering, which is also treated analytically.
To accomplish our objective, we structure this paper as follows: In Section II, we introduce the spacetime under study and develop a Lagrangian formalism, leading to the equations of motion for the test particles. In Section III, we initiate our main study by providing a general classification of the possible orbits based on the radial profile of the effective potential experienced by the particles. In this section, we provide a precise derivation of the planetary orbits with different eccentricities, also assessed by observational tests regarding the S-star orbits at the center of the Milky Way. In Section IV, we continue our study by investigating the unbound orbits, consisting of deflecting trajectories. These trajectories are either deflected beyond the black hole or inexorably fall into the event horizon. The former is termed gravitational Rutherford scattering in Refs. Villanueva and Olivares (2015); Fathi et al. (2021a). To complete this investigation, in Section V, we present a fully analytical study of the purely radial trajectories. We conclude in Section VI. Throughout the paper, we work in the geometrized system of units where $G=c=1$, we adopt the sign convention ($+$)(-++\,+)( - + + + ), and wherever primes appear, they denote differentiations with respect to the radial coordinate.
@@ II.1 Magnetization of Schwarzschild-like black holes
$F_{\theta\phi}=\frac{B}{2}r^{2}\sin{2\theta}.$ (3b)
@@ II.2 Magnetized particle dynamics
with $\mathcal{S}$ denoting the Jacobi action, and motion restricted to the equatorial plane, the action can be expressed as ${\cal S}=-Et+L\phi+{\cal S}_{r}(r)$, and in this context, we have $m=1$. Therefore, the radial motion can be easily described in terms of the following equation:
@@
$V_{\mathrm{eff}}(r;l)=f(r)\left(\left[1-{\beta\mathcal{F}(r,\pi/2)}\right]^{2}+\frac{l^{2}}{r^{2}}\right).$ (13)
is the effective potential felt by the test particles.
@@
In Fig. 1, we illustrate how the radial dependence of the effective potential varies with the specific angular momentum (left panel) and the magnetic interaction parameter $\beta$ (middle panel). The maximum of the effective potential denotes instability. Furthermore, these plots indicate that increasing the specific angular momentum $l$ of the magnetized particles and the magnetic interaction parameter $\beta$ leads to an increase in the stable circular orbits (at $r_{\text{min}}$) of the test particles with magnetic dipole momentum. Conversely, an increase in the specific angular momentum of the magnetized particle results in a decrease in the unstable circular orbits (at $r_{\text{max}}$) of the particles with magnetic dipole moment. Moreover, escalating the magnetic coupling parameter $\beta$ leads to an increase in the unstable circular orbits of the magnetized particles. Additionally, Fig. 1 presents the radial profile of the effective potential with the specific angular momentum $l_{\text{ISCO}}$ of the magnetized particles in the Innermost Stable Circular Orbit (ISCO) for a given value of $\beta=0.1$. Consequently, Fig. 1 indicates that for specific angular momentum values $l>l_{\text{ISCO}}$, there are two extreme points in the effective potential, namely $r_{\text{max}}$ corresponding to unstable circular orbits and $r_{\text{min}}$ corresponding to stable circular orbits. For the specific angular momentum value $l=l_{\text{ISCO}}$, there is only one extreme point $r_{\text{min}}$, corresponding to the stable circular orbit. However, for specific angular momentum values $l<l_{\text{ISCO}}$, there are no extreme points in the effective potential, indicating the absence of any circular orbits in this case. Additionally, insights into the orbits of magnetized test particles can be obtained from the effective potential. In the right panel of Fig. 1, we depict the turning points $r_{s}$ and $r_{l}$ corresponding to the deflecting trajectories. Furthermore, concerning the planetary bound orbits, we have also denoted the turning points $r_{p}$ (periapsis), the smallest orbital separation, and $r_{a}$ (apoapsis), the largest orbital separation, of particles possessing magnetic dipole moment orbiting a Schwarzschild Black Hole within magnetized fields. We will later explore that test particles establish bound orbits if their specific energy $\mathcal{E}$ is less than $\sqrt{1-\beta}$ and unbound orbits if $\mathcal{E}_{u}>\mathcal{E}\geq\sqrt{1-\beta}$, where $\mathcal{E}_{u}^{2}=V_{\text{eff}}(r_{u})$ represents the energy corresponding to the unstable circular orbits at $r_{u}=r_{\text{max}}$.
@@
In the forthcoming sections of the paper, we will employ this equation to analytically analyse the bound and unbound orbits of magnetized test particles in the spacetime under consideration.
@@ III The bound orbits
Following the same procedure as in Ref. Lim and Yeo (2024) to determine the ($l,\mathcal{E}$) parameter space for circular orbits, in this section we will investigate how this space varies with the magnetic coupling constant $\beta$. Subsequently, by employing the conditions $\mathcal{E}^{2}=V_{\text{eff}}$ and $V_{\text{eff}}^{\prime}=0$, we can ascertain the specific energy $\mathcal{E}$ and specific angular momentum $l$ of the magnetized particles as
@@
corresponding to circular orbits at the radius $r=r_{0}$. It is well-known that to identify stable circular orbits, one should examine the second derivative of the effective potential, namely $V_{\mathrm{eff}}^{\prime\prime}>0$ indicates stable circular orbits, while $V_{\mathrm{eff}}^{\prime\prime}<0$ indicates unstable circular orbits. However, our determination of stable and unstable circular orbits depends on $\beta$. Therefore, in Table 1, we present the values of the critical radii of the unstable circular orbits. It is evident that $V_{\mathrm{eff}}^{\prime\prime}>0$ for $r_{0}>r_{\text{max}}$, corresponding to stable orbits, and $V_{\mathrm{eff}}^{\prime\prime}<0$ for $r_{\text{min}}<r_{0}<r_{\text{max}}$, corresponding to unstable orbits. Surely, $r_{0}=r_{\text{min}}$ corresponds to the ISCO. Therefore, from Table 1, one can easily conclude that increasing the magnetic coupling constant $\beta$ leads to an increase in the value of the ISCO radius. To provide more precise insight, in Fig. 2, we
@@ 0.0 $3M$ $6M$
illustrate how the ISCO radius $r_{\text{ISCO}}$, the specific energy $\mathcal{E}_{\text{ISCO}}$, and the specific angular momentum $l_{\text{ISCO}}$ of the magnetized particle in the ISCO depend on the magnetic coupling constant $\beta$. From Fig. 2, one can readily conclude that increasing the magnetic coupling parameter $\beta$ results in a decrease in the specific energy $\mathcal{E}_{\text{ISCO}}$ and the specific angular momentum $l_{\text{ISCO}}$ of the magnetized particles in the ISCO. In contrast, increasing the magnetic interaction term $\beta$ leads to an increase in the value of the ISCO radius $r_{\text{ISCO}}$ of the particles with the magnetic dipole moment. Additionally, in Fig. 3, we have plotted the domain (for the quantities $\mathcal{E}$, $l/M$, and the line representing the non-relativistic limit, i.e. $l/M\to\infty$), within which, magnetized particles can possess circular orbits.
@@ III.1 Analytical solution for the angular trajectories
Now, to obtain a general analytical solution for the angular motion, we assume that the magnetized particles commence their motion from $u=u_{1}$ and $\phi=0$. Then, Eq. (18) can be expressed as
@@
where $\theta_{0}\equiv\theta_{0}(\phi)=\sqrt{a(1+\beta)(u_{3}-u_{1})}{\phi}/{2}$. In the above relation, $\operatorname{\mathrm{sn}}(\theta,k)^{}$ [^1] The well-known incomplete elliptic integral of the first kind is given by the general form Byrd and Friedman (1971) $\theta=F(\phi,k)=\int_{0}^{\phi}\frac{dt}{\sqrt{1-k^{2}\sin^{2}{t}}},$ where $\phi=F^{-1}(\theta,k)=\mathrm{am}(\theta,k)$ is the Jacobi amplitude. This way, the Jacobi elliptic sine function is defined as $\sin{\phi}=\sin{\big{(}\mathrm{am}(\theta,k)\big{)}=\operatorname{\mathrm{sn}}(\theta,k)}.$ denotes the Jacobi elliptic sine function, with argument $\theta$ and modulus $k$.

### v2

@@ Abstract
This study provides a comprehensive analytical investigation of the bound and unbound motion of magnetized particles orbiting a Schwarzschild black hole immersed in an external asymptotically uniform magnetic field, which includes all conceivable types of bounded and unbounded orbits. In particular, for planetary orbits, we perform a comparative analysis of our findings with the observed position of the S2 star carrying magnetic dipole moment around Sagittarius A* (Sgr A*). We found maximum and minimum values for the parameter of magnetic interaction between the magnetic dipole of the star and the external magnetic field, as well as the energy and angular momentum of the S2 star. As a result, we obtain estimations of the magnetic dipole of the star in order of $10^{6}\rm\ G\cdot cm^{3}$. Additionally, we explore deflecting trajectories akin to gravitational Rutherford scattering. In obtaining the solutions for the orbital equations, we articulate the elliptic integrals and Jacobi elliptic functions, and illustrative figures and simulations augment our study.
@@ I Introduction
The investigation of free-falling objects in gravitational fields links geometry and gravity through the general theory of relativity, where the curvature of spacetime plays the principal role. This fundamental understanding gained popularity early once it was recognized that planets and light follow geodesics. The outstanding discovery, right after the Schwarzschild solution proposition Schwarzschild (1916), was the verification of light deflection during the 1919 solar eclipse expedition Dyson et al. (1920). The precise evaluation of the anomalous precession in the perihelion of Mercury Clemence (1947), which constitute the primary standard tests of general relativity. However, the complex nature of the partial differential equations governing particle dynamics in curved spacetimes necessitates simplified or numerical solutions to compare theoretical results with observational tests.
@@
Analytical expressions have several advantages since they serve as benchmarks for numerical methods and, hence, they facilitate comprehensive parameter space exploration for predicting astrophysical observables. Efforts to find exact analytical solutions for geodesic equations of particles have intensified since Hagihara’s studies in 1931 Hagihara (1930), followed by works by Darwin, Mielnik, and Plebański Darwin (1959, 1961); Mielnik and Plebański (1962). These methods, which are based on the theories of elliptic functions and modular forms, were studied by nominated nineteenth-century mathematicians such as Jacobi Jacobi (2013), Abel Abel , Riemann Riemann (1857, 1866), and Weierstrass Weierstrass (1854) (see also Ref. Baker (1995) for a complete textbook review on these discoveries). Particularly noteworthy is the application of modular forms in solving (hyper-)elliptic integrals involved in particle trajectory studies, garnering significant attention in recent decades. These methods, rooted in the theories of elliptic functions and modular forms, have been the focus of numerous investigations analyzing trajectory curves in black hole spacetimes inferred from general relativity and its extensions (see, for example, Refs. Kraniotis and Whitehouse (2002, 2003); Kraniotis (2004, 2005); Cruz et al. (2005); Kraniotis (2007); Hackmann and Lämmerzahl (2008a, b); Hackmann et al. (2009, 2010); Olivares et al. (2011); Kraniotis (2011); Cruz et al. (2013); Villanueva et al. (2013); Kraniotis (2014); Soroushfar et al. (2015); Villanueva and Olivares (2015); Soroushfar et al. (2016); Hoseini et al. (2016, 2017); Fathi et al. (2020a, b, 2021a); González et al. (2021); Fathi et al. (2021b); Kraniotis (2021); Fathi et al. (2022); Soroushfar and Afrooz (2022); Battista and Esposito (2022); Fathi et al. (2023); Fathi (2023)).
@@
Building upon this scientific foundation, contemporary research continues to explore the properties and dynamics of particles falling in gravitational fields in diverse astrophysical contexts. bound and unbound particle orbits, in particular, represent fundamental aspects of motion that offer valuable insights into the structure and evolution of the universe. Bound orbits, characterized by closed trajectories around central masses, are found extensively in celestial systems, ranging from planetary orbits in the solar system to the motion of stars in galaxies Perek (1962); Barbieri and Gratton (2002). These orbits contribute to astrophysical objects’ stability and long-term dynamics, and provide essential constraints on theoretical models of gravitational dynamics and cosmological evolution.
@@
Given the importance of this subject, our study aims to investigate the dynamics of magnetized particles affected by the magnetic field of an astrophysical object, in this case, a magnetized Schwarzschild black hole. In this study, we analyze the dynamics and motion of particles possessing magnetic dipole moments analytically as they move in the exterior geometry of a Schwarzschild black hole with an associated magnetic field. We utilize elliptic integrals to express exact analytical solutions to the equations of motion governing the behavior of these particles. Particularly, the deflecting trajectories are akin to a type of gravitational Rutherford scattering, which is also treated analytically.
To accomplish our objective, we structure this paper as follows: In Section II, we introduce the spacetime under study and develop a Lagrangian formalism, leading to the equations of motion for the test particles. In Section III, we initiate our main study by providing a general classification of the possible orbits based on the radial profile of the effective potential experienced by the particles. In this section, we provide a precise derivation of the planetary orbits with different eccentricities, also assessed by observational tests regarding the S-star orbits at the center of the Milky Way. In Section IV, we continue our study by investigating the unbound orbits consisting of deflecting trajectories. These trajectories are deflected beyond the black hole or inevitably fall into the event horizon. The former is termed gravitational Rutherford scattering in Refs. Villanueva and Olivares (2015); Fathi et al. (2021a). To complete this investigation, in Section V, we present a fully analytical study of the purely radial trajectories. We conclude in Section VI. Throughout the paper, we work in the geometrized system of units where $G=c=1$, we adopt the sign convention ($+$)(-++\,+)( - + + + ), and wherever primes appear, they denote differentiations for the radial coordinate.
@@ II.1 Magnetization of Schwarzschild-like black holes
$F_{\theta\phi}=Br^{2}\sin{\theta}\cos\theta.$ (3b)
@@ II.2 Magnetized particle dynamics
with $\mathcal{S}$ denoting the Jacobi action, and motion restricted to the equatorial plane, the action can be expressed as ${\cal S}=-Et+L\phi+{\cal S}_{r}(r)$, and in this context. Therefore, the radial motion can be easily described in terms of the following equation:
@@
$V_{\mathrm{eff}}(r;l)=f(r)\left(\left[1-{\beta\mathcal{F}(r,\pi/2)}\right]^{2}+\frac{l^{2}}{r^{2}}\right),$ (13)
do the test particles feel the effective potential, which takes the following form in the weak electromagnetic interaction limit (at $\beta\ll 1$), considering the linear order of $\beta$:
$V_{\mathrm{eff}}(r;l)=f(r)\left(1-2{\beta\mathcal{F}(r,\pi/2)}+\frac{l^{2}}{r^{2}}\right).$ (14)
@@
In Fig. 1, we illustrate how the radial dependence of the effective potential varies with the specific angular momentum (left panel) and the magnetic interaction parameter $\beta$ (middle panel). The maximum of the effective potential denotes instability. Furthermore, these plots indicate that increasing the specific angular momentum $l$ of the magnetized particles and the magnetic interaction parameter $\beta$ leads to an increase in the stable circular orbits (at $r_{\text{min}}$) of the test particles with magnetic dipole momentum. Conversely, an increase in the specific angular momentum of the magnetized particle results in a decrease in the unstable circular orbits (at $r_{\text{max}}$) of the particles with magnetic dipole moment. Moreover, escalating the magnetic coupling parameter $\beta$ increases the unstable circular orbits of the magnetized particles. Additionally, Fig. 1 presents the radial profile of the effective potential with the specific angular momentum $l_{\text{ISCO}}$ of the magnetized particles in the Innermost Stable Circular Orbit (ISCO) for a given value of $\beta=0.1$. Consequently, Fig. 1 indicates that for specific angular momentum values $l>l_{\text{ISCO}}$, there are two extreme points in the effective potential, namely $r_{\text{max}}$ corresponding to unstable circular orbits and $r_{\text{min}}$ corresponding to stable circular orbits. For the specific angular momentum value $l=l_{\text{ISCO}}$, there is only one extreme point $r_{\text{min}}$, corresponding to the stable circular orbit. However, for specific angular momentum values $l<l_{\text{ISCO}}$, there are no extreme points in the effective potential, indicating the absence of any circular orbits in this case. Additionally, insights into the orbits of magnetized test particles can be obtained from the effective potential. In the right panel of Fig. 1, we depict the turning points $r_{s}$ and $r_{l}$ corresponding to the deflecting trajectories. Furthermore, concerning the planetary bound orbits, we have also denoted the turning points $r_{p}$ (periapsis), the smallest orbital separation, and $r_{a}$ (apoapsis), the most significant orbital separation, of particles possessing magnetic dipole moment orbiting a Schwarzschild Black Hole within magnetized fields. We will later explore that test particles establish bound orbits if their specific energy $\mathcal{E}$ is less than $\sqrt{1-\beta}$ and unbound orbits if $\mathcal{E}_{u}>\mathcal{E}\geq\sqrt{1-\beta}$, where $\mathcal{E}_{u}^{2}=V_{\text{eff}}(r_{u})$ represents the energy corresponding to the unstable circular orbits at $r_{u}=r_{\text{max}}$.
@@
By simplifying Eq. (13) to the first order in $\beta$, we obtain
$\left(\frac{dr}{d\phi}\right)^{2}=\frac{r^{4}}{l^{2}}\Biggl{\{}\mathcal{E}^{2}\left(1+2\beta\sqrt{f(r)}\right)\\ -f(r)\left[\beta\sqrt{f(r)}\left(1+\frac{2l^{2}}{r^{2}}\right)+1+\frac{l^{2}}{r^{2}}\right]\Biggr{\}}.$ (16)
In the forthcoming paper sections, we will employ this equation to analyze the bound and unbound orbits of magnetized test particles in the spacetime under consideration.
@@ III The bound orbits
Following the same procedure as in Ref. Lim and Yeo (2024) to determine the ($l,\mathcal{E}$) parameter space for circular orbits, in this section, we investigate how this space varies with the magnetic coupling constant $\beta$. Subsequently, by employing the conditions $\mathcal{E}^{2}=V_{\text{eff}}$ and $V_{\text{eff}}^{\prime}=0$, we can ascertain the specific energy $\mathcal{E}$ and specific angular momentum $l$ of the magnetized particles as
@@
corresponding to circular orbits at the radius $r=r_{0}$. It is well-known that to identify stable circular orbits, one should examine the second derivative of the effective potential, namely $V_{\mathrm{eff}}^{\prime\prime}>0$ indicates stable circular orbits. In contrast, $V_{\mathrm{eff}}^{\prime\prime}<0$ indicates unstable circular orbits. However, our determination of stable and unstable circular orbits depends on $\beta$. Therefore, in Table 1, we present the values of the critical radii of the unstable circular orbits. It is evident that $V_{\mathrm{eff}}^{\prime\prime}>0$ for $r_{0}>r_{\text{max}}$, corresponding to stable orbits, and $V_{\mathrm{eff}}^{\prime\prime}<0$ for $r_{\text{min}}<r_{0}<r_{\text{max}}$, corresponding to unstable orbits. Surely, $r_{0}=r_{\text{min}}$ corresponds to the ISCO. Therefore, from Table 1, one can easily conclude that increasing the magnetic coupling constant $\beta$ leads to an increase in the value of the ISCO radius. To provide more precise insight, in Fig. 2, we
@@ 0.0 $3M$ $6M$
illustrate how the ISCO radius $r_{\text{ISCO}}$, the specific energy $\mathcal{E}_{\text{ISCO}}$, and the specific angular momentum $l_{\text{ISCO}}$ of the magnetized particle in the ISCO depend on the magnetic coupling constant $\beta$. From Fig. 2, one can readily conclude that increasing the magnetic coupling parameter $\beta$ results in a decrease in the specific energy $\mathcal{E}_{\text{ISCO}}$ and the specific angular momentum $l_{\text{ISCO}}$ of the magnetized particles in the ISCO. In contrast, increasing the magnetic interaction term $\beta$ leads to a rise in the value of the ISCO radius $r_{\text{ISCO}}$ of the particles with the magnetic dipole moment. Additionally, in Fig. 3, we have plotted the domain (for the quantities $\mathcal{E}$, $l/M$, and the line representing the non-relativistic limit, i.e. $l/M\to\infty$), within which, magnetized particles can possess circular orbits.
@@ III.1 Analytical solution for the angular trajectories
In the weak interaction limit, Eq. (20) can be recast as
$\left(\frac{du}{d\phi}\right)^{2}=a_{0}u^{3}-b_{0}u^{2}+c_{0}u+d_{0},$ (22)
in which
$a_{0}=2M(1+3\beta),$ (23a)
$b_{0}=1+2\beta\left(1+\frac{M^{2}}{l^{2}}\right),$ (23b)
$c_{0}=\frac{M}{l^{2}}\bigl{[}2+\beta(3-2\mathcal{E}^{2})\bigr{]},$ (23c)
$d_{0}=\frac{\beta(2\mathcal{E}^{2}-1)+(\mathcal{E}^{2}-1)}{l^{2}}.$ (23d)
@@
To obtain a general analytical solution for the angular motion, we assume that the magnetized particles commence their motion from $u=u_{1}$ and $\phi=0$. Then, Eq. (20) can be expressed as
@@
where $\theta_{0}\equiv\theta_{0}(\phi)=\sqrt{a(1+\beta)(u_{3}-u_{1})}\,{\phi}/{2}$. In the above relation, $\operatorname{\mathrm{sn}}(\psi,k)$ denotes the Jacobi elliptic sine function with argument $\psi$ and modulus $k$, and is defined as $\operatorname{\mathrm{sn}}(\psi,k)=\sin\xi$, for which
$\psi=F(\xi,k)=\int_{0}^{\xi}\frac{dt}{\sqrt{1-k^{2}\sin^{2}t}}.$ (28)